# Hardware Acceleration for Neural Networks: A Comprehensive Survey

### v1

@@ 3.4.1. Inference on FPGAs
Inference on FPGAs has a long history of efficient CNN and BNN implementations using quantization, folding, and pipeline parallelism (Umuroglu et al., 2017; Ma et al., 2018). Designs typically trade flexibility for efficiency by fixing layer parameters and building a streaming pipeline that reuses on-chip buffers and minimizes off-chip traffic. When the model changes frequently or includes operators outside the accelerator’s coverage, the benefits can diminish due to host fallback and data marshaling overhead (Nurvitadhi et al., 2017; Venieris and Bouganis, 2017).
Recent work also targets Transformer inference by accelerating attention, softmax, and feed-forward layers with structured sparsity patterns and custom compute engines (Zhang et al., 2024a). Attention-heavy workloads stress memory bandwidth and require efficient buffering strategies for QKV projections and intermediate activations; they also benefit from kernel fusion to reduce intermediate writes. Compared to GPUs, FPGAs can deliver strong latency determinism and energy efficiency, but performance depends heavily on external memory bandwidth, operator coverage, and toolchain quality (HLS scheduling, memory partitioning, and DSP utilization) (Nurvitadhi et al., 2017).
@@ 3.5. LLM-serving accelerators (LPUs and related designs)
Recently, specialized accelerators have emerged targeting the inference-serving regime of large language models. These systems often emphasize predictable execution, low-latency token generation, and high throughput under real-time constraints, with architecture and software co-designed around attention-centric workloads and memory management (Groq, 2024; Kwon et al., 2023). This deployment regime differs from classical batch inference: it combines strict latency targets with fluctuating concurrency, diverse prompt lengths, and complex scheduling policies for multi-tenant serving. As a result, accelerator design must address not only peak GEMM throughput but also memory capacity, memory bandwidth, and the overheads of dynamic batching and KV-cache allocation (Vaswani et al., 2017; Kwon et al., 2023).
@@ 3.5.1. Inference on LPUs
Figure 13. Inference on LPUs
@@ 3.7. Neuromorphic accelerators
Figure 14. Neuromorphic accelerator architecture
@@ 4. Accelerator Approaches
Figure 15. accelerator architectures for different neural networks
@@ 4.1. ANN acceleration (MLP and fully connected networks)
Figure 16. ANN architecture
@@ 4.1.4. ANNs on FPGAs
On FPGAs, ANN acceleration often uses spatial pipelines: a GEMM is implemented as a tiled matrix-multiply engine with customized precision, and data are streamed through compute stages to avoid intermediate DRAM writes (Nurvitadhi et al., 2017; Venieris and Bouganis, 2017). Rather than launching kernels, the design builds a fixed datapath where each cycle advances partial sums through a pipeline, enabling deterministic latency when the workload shape is fixed. Tiling controls how much parallelism is unrolled in hardware and how operands are buffered on-chip (Williams et al., 2009).
@@
Performance is frequently bounded by external memory bandwidth and by the overhead of moving tensors across the host–FPGA interface in system deployments (Nurvitadhi et al., 2017). Consequently, many effective FPGA ANN designs try to keep weights resident on-chip (when feasible), stream activations, and fuse adjacent elementwise operations into the datapath so that the pipeline remains compute-fed rather than I/O-limited (Williams et al., 2009; Venieris and Bouganis, 2017).
@@ 4.2. CNN acceleration
Figure 17. CNN architecture
Compared to generic dense layers, CNN accelerators can achieve high energy efficiency because reuse is predictable and amenable to spatial architectures and streaming pipelines (e.g., ASIC and FPGA designs); Figure 17 summarizes a representative CNN structure (Chen et al., 2016; Umuroglu et al., 2017; Venieris and Bouganis, 2017).
@@ 4.2.4. CNNs on FPGAs
FPGA CNN accelerators often use streaming dataflows where feature maps flow through a pipeline of convolution/activation stages, avoiding intermediate off-chip traffic and providing deterministic latency (Venieris and Bouganis, 2017). A common pattern is to use line buffers and sliding-window datapaths that reuse input pixels across neighboring convolution windows, matching convolution’s spatial locality. The pipeline can be “folded” to trade throughput for area by time-multiplexing a smaller compute engine across channels or layers, while still maintaining a streaming schedule (Venieris and Bouganis, 2017; Williams et al., 2009).
@@
Practically, the design must balance DSP utilization, BRAM partitioning, and external memory bandwidth; otherwise the pipeline stalls on memory rather than compute (Nurvitadhi et al., 2017; Williams et al., 2009). System-level bottlenecks such as host I/O (PCIe), buffering, and reconfiguration overhead also affect end-to-end performance in deployment settings, motivating designs that stream directly from input sources and fuse as many layers as feasible into a single datapath (Nurvitadhi et al., 2017; Venieris and Bouganis, 2017).
@@ 4.3. RNN acceleration
  • Gate fusion: fuse multiple gate computations (e.g., in LSTM/GRU-style cells) into fewer kernels to reduce memory traffic and kernel-launch overhead, improving latency at small batch sizes (Chetlur et al., 2014; Tillet et al., 2019).
@@
Figure 18. RNN architecture
@@ 4.3.4. RNNs on FPGAs
However, long sequences can stress on-chip buffering, and performance is often bounded by external memory bandwidth or host I/O if the pipeline cannot keep data resident (Williams et al., 2009; Venieris and Bouganis, 2017). In system deployments, additional overhead can come from host-device transfers and from how sequences are batched/packed; therefore, practical FPGA RNN acceleration often co-designs the data ingress/egress pipeline with the compute datapath so that the recurrent loop remains streaming rather than I/O-bound (Nurvitadhi et al., 2017; Williams et al., 2009).
@@ 4.4.1. Different Transformer and LLM model variants
  • Encoder-only Transformers: used for representation learning (e.g., classification/embedding). They process the full sequence in parallel, so throughput is driven by batched matmuls and attention over the full context (e.g., BERT-style encoders) (Vaswani et al., 2017; Devlin et al., 2019).
@@
  • Decoder-only LLMs: used for autoregressive generation (e.g., GPT-style and LLaMA-style). Inference alternates between a prefill phase (processing the prompt) and a decode phase (generating tokens). Decode is often bandwidth-limited because each token requires reading/writing KV-cache and executing relatively small GEMMs (Brown et al., 2020; Touvron et al., 2023; Kwon et al., 2023; Williams et al., 2009).
@@ 4.4.2. Key kernels and bottlenecks: attention, KV-cache, and fusion
Figure 19. Kernels and bottlenecks for accelerating LLM
@@
Recent acceleration strategies for Transformer/LLM inference increasingly target decode efficiency and memory traffic per token. Examples include faster IO-aware attention implementations beyond the original FlashAttention design (Dao et al., 2022; Dao, 2023), attention variants that reduce KV-cache bandwidth (e.g., grouped-query attention, which reduces the number of KV heads relative to Q heads) (Ainslie et al., 2023), and serving-time techniques such as speculative decoding that trade extra parallel work for lower end-to-end latency (Leviathan et al., 2023). Post-training quantization has also evolved into practical, high-accuracy pipelines for LLMs—including SmoothQuant, GPTQ, and activation-aware weight quantization (AWQ)—which reduce bandwidth and enable lower-precision kernels when supported by hardware and libraries (Xiao et al., 2022; Frantar and Alistarh, 2022; Lin et al., 2023; Jacob et al., 2018).
@@ 4.4.3. Transformers/LLMs on GPUs
When MoE is used, GPU efficiency depends on routing implementation, expert parallelism, and load balance; otherwise, conditional execution can reduce utilization and increase tail latency even when compute capacity is high (Shazeer et al., 2017; Gale et al., 2019). Overall, GPUs remain the default platform because the software ecosystem (libraries, compilers, serving runtimes) evolves rapidly with model variants and can incorporate new attention kernels and scheduling policies without changing hardware (Chen et al., 2018; Tillet et al., 2019).
@@ 4.4.4. Transformers/LLMs on TPUs/NPUs
For large-scale training, TPU-class systems demonstrate that system co-design (HBM bandwidth, on-chip SRAM, and high-bandwidth interconnect) matters as much as per-chip compute. Scaling LLM training requires sharding and collective communication that interact with compilation decisions, and overall efficiency depends on overlap between compute and communication as well as on memory footprint (activations, optimizer state) (Jouppi et al., 2021; Micikevicius et al., 2018). Reduced precision (BF16/FP16) is widely used to increase throughput and manage memory traffic (Micikevicius et al., 2018).
@@ 4.4.6. Transformers/LLMs on FPGAs
FPGA acceleration(Fig.20) for Transformers focuses on building streaming pipelines for projections/MLP and (when feasible) attention/softmax, using custom precision and on-chip buffering to reduce off-chip traffic (Zhang et al., 2024a; Nurvitadhi et al., 2017). The FPGA approach mirrors classical CNN FPGA acceleration: exploit spatial pipelining and determinism, tailor datapath precision, and fuse adjacent stages so that intermediate tensors do not spill to DRAM (Venieris and Bouganis, 2017; Williams et al., 2009).
Figure 20. FPGA accelerating pipeline
@@
System integration is a major determinant of end-to-end benefit. Host I/O (e.g., PCIe), data marshaling, and batching policies can dominate latency unless the serving pipeline is engineered so that the FPGA remains compute-fed (Nurvitadhi et al., 2017). Consequently, FPGA designs often target fixed model shapes and low-latency regimes where deterministic pipelines and tight I/O integration offset lower peak compute, while toolchain support and memory-system design remain central to achieving robust speedups (Venieris and Bouganis, 2017; Nurvitadhi et al., 2017).
@@
Beyond mainstream GPU/TPU deployments, the 2023–2025 landscape also includes a growing set of alternative datacenter accelerators (e.g., AMD Instinct MI300-class systems, Intel Gaudi-class accelerators, and AWS Trainium-class instances) (Wikipedia contributors, 2025a, c, b). On the research side, there is active work on transformer-specific accelerators and system co-design: dynamic sparsity/activation pruning in transformer blocks (e.g., AccelTran) (Wang et al., 2023), 3D heterogeneous manycore organizations for end-to-end transformer execution (e.g., HeTraX) (Zhang et al., 2024b), FPGA toolflow-driven LLM inference (e.g., HLS-based transformer implementations) (Li et al., 2024), and heterogeneous processing-in-memory (PIM) approaches that couple SRAM/HBM compute to reduce bandwidth bottlenecks (Kim et al., 2025). These efforts reinforce the same overarching lesson: as models scale and contexts grow, end-to-end speedups come from jointly optimizing kernels, memory systems, and runtimes—not from compute throughput alone (Williams et al., 2009; Kwon et al., 2023).
@@ 6. Conclusion
Neural-network acceleration has evolved from improving peak arithmetic throughput to optimizing end-to-end execution under memory, communication, and systems constraints. Across GPUs, TPUs/NPUs, FPGAs, and ASIC accelerators, the most consistent performance and energy gains come from reducing data movement through locality-aware dataflows, larger and better-managed memory hierarchies, reduced precision, and operator fusion—often enabled by compilers and runtimes that map high-level models to hardware efficiently (Williams et al., 2009; Hennessy and Patterson, 2019).
Looking forward, the deployment regime of large language models highlights several open challenges. Long-context and high-concurrency serving make KV-cache capacity and bandwidth first-class bottlenecks, requiring IO-aware attention kernels and robust runtime memory management (Dao et al., 2022; Kwon et al., 2023). Model heterogeneity (e.g., MoE and tool-augmented inference) increases conditional execution and load imbalance, pushing accelerator stacks toward richer scheduling support without sacrificing efficiency (Shazeer et al., 2017; Gale et al., 2019). Finally, fair comparison remains difficult in a rapidly changing ecosystem; reproducible evaluation should report end-to-end metrics and the full software stack, since kernels, compilation, and serving policy can dominate realized performance (Mattson et al., 2020; Williams et al., 2009).

### v3

@@ 3.4.1. Inference on FPGAs
Figure 12 sketches a representative FPGA-based acceleration setup used to discuss streaming datapaths, buffering, and I/O constraints. Inference on FPGAs has a long history of efficient CNN and Binarized Neural Network (BNN) implementations using quantization, folding, and pipeline parallelism (Umuroglu et al., 2017; Ma et al., 2018). Designs typically trade flexibility for efficiency by fixing layer parameters and building a streaming pipeline that reuses on-chip buffers and minimizes off-chip traffic. When the model changes frequently or includes operators outside the accelerator’s coverage, the benefits can diminish due to host fallback and data marshaling overhead (Nurvitadhi et al., 2017; Venieris and Bouganis, 2016).
Recent work also targets Transformer inference by accelerating attention, softmax, and feed-forward layers with structured sparsity patterns and custom compute engines (Zhang et al., 2024a). Attention-heavy workloads stress memory bandwidth and require efficient buffering strategies for QKV projections and intermediate activations; they also benefit from kernel fusion to reduce intermediate writes. Compared to GPUs, FPGAs can deliver strong latency determinism and energy efficiency, but performance depends heavily on external memory bandwidth, operator coverage, and toolchain quality (HLS scheduling, memory partitioning, and DSP utilization) (Nurvitadhi et al., 2017). Recent work has also demonstrated the efficacy of FPGAs for accelerating Digital Twin learning and model recovery in mission-critical edge applications (Xu et al., 2025e, c, b).
@@ 3.5. LLM-serving accelerators (LPUs and related designs)
Recently, specialized accelerators have emerged targeting the inference-serving regime of large language models. These systems often emphasize predictable execution, low-latency token generation, and high throughput under real-time constraints, with architecture and software co-designed around attention-centric workloads and memory management (Groq, 2024; Kwon et al., 2023). Recent surveys highlight that achieving “faster and lighter” LLMs requires a combination of algorithmic compression (quantization, pruning) and system-level optimization, reinforcing that hardware cannot be designed in isolation from the serving stack (Chavan et al., 2024). This deployment regime differs from classical batch inference: it combines strict latency targets with fluctuating concurrency, diverse prompt lengths, and complex scheduling policies for multi-tenant serving. As a result, accelerator design must address not only peak GEMM throughput but also memory capacity, memory bandwidth, and the overheads of dynamic batching and KV-cache allocation (Vaswani et al., 2017; Kwon et al., 2023).
@@ 3.5.1. Inference on LPUs
Figure 13. Inference on LPU-style accelerators, focusing on deterministic low-latency token generation for LLMs by managing control flow in hardware and optimizing memory access patterns for the decode phase.
@@ 3.7. Neuromorphic accelerators
Figure 14. Neuromorphic accelerator architecture emphasizing event-driven processing, distributed local memory, and asynchronous communication to achieve high energy efficiency for sparse workloads.
@@ 4. Accelerator Approaches
Figure 15. Taxonomy of accelerator architectures tailored for different neural network types (ANNs, CNNs, RNNs, Transformers), showing how workload characteristics drive specific hardware design choices.
@@ 4.1. ANN acceleration (MLP and fully connected networks)
Figure 16. Standard Artificial Neural Network (ANN) structure used as a reference for acceleration discussions, featuring dense fully connected layers that benefit from matrix-multiplication optimization.
@@ 4.1.4. ANNs on FPGAs
On FPGAs, ANN acceleration often uses spatial pipelines: a GEMM is implemented as a tiled matrix-multiply engine with customized precision, and data are streamed through compute stages to avoid intermediate DRAM writes (Nurvitadhi et al., 2017; Venieris and Bouganis, 2016). Rather than launching kernels, the design builds a fixed datapath where each cycle advances partial sums through a pipeline, enabling deterministic latency when the workload shape is fixed. Tiling controls how much parallelism is unrolled in hardware and how operands are buffered on-chip (Williams et al., 2009).
@@
Performance is frequently bounded by external memory bandwidth and by the overhead of moving tensors across the host–FPGA interface in system deployments (Nurvitadhi et al., 2017). Consequently, many effective FPGA ANN designs try to keep weights resident on-chip (when feasible), stream activations, and fuse adjacent elementwise operations into the datapath so that the pipeline remains compute-fed rather than I/O-limited (Williams et al., 2009; Venieris and Bouganis, 2016).
@@ 4.2. CNN acceleration
Figure 17. Convolutional Neural Network (CNN) structure highlighting spatial locality and weight reuse, which are exploited by hardware through specialized dataflows and buffering strategies.
Compared to generic dense layers, CNN accelerators can achieve high energy efficiency because reuse is predictable and amenable to spatial architectures and streaming pipelines (e.g., ASIC and FPGA designs); Figure 17 summarizes a representative CNN structure (Chen et al., 2016; Umuroglu et al., 2017; Venieris and Bouganis, 2016).
@@ 4.2.4. CNNs on FPGAs
FPGA CNN accelerators often use streaming dataflows where feature maps flow through a pipeline of convolution/activation stages, avoiding intermediate off-chip traffic and providing deterministic latency (Venieris and Bouganis, 2016; Yan et al., 2024). A common pattern is to use line buffers and sliding-window datapaths that reuse input pixels across neighboring convolution windows, matching convolution’s spatial locality. The pipeline can be “folded” to trade throughput for area by time-multiplexing a smaller compute engine across channels or layers, while still maintaining a streaming schedule (Venieris and Bouganis, 2016; Williams et al., 2009).
@@
Practically, the design must balance DSP utilization, BRAM partitioning, and external memory bandwidth; otherwise the pipeline stalls on memory rather than compute (Nurvitadhi et al., 2017; Williams et al., 2009). System-level bottlenecks such as host I/O (PCIe), buffering, and reconfiguration overhead also affect end-to-end performance in deployment settings, motivating designs that stream directly from input sources and fuse as many layers as feasible into a single datapath (Nurvitadhi et al., 2017; Venieris and Bouganis, 2016).
@@ 4.3. RNN acceleration
  • Gate fusion: fuse multiple gate computations (e.g., in Long Short-Term Memory (LSTM)/Gated Recurrent Unit (GRU)-style cells) into fewer kernels to reduce memory traffic and kernel-launch overhead, improving latency at small batch sizes (Chetlur et al., 2014; Tillet et al., 2019).
@@
Figure 18. Recurrent Neural Network (RNN) structure showing temporal dependencies and feedback loops, which present challenges for parallelization and require specialized state-buffering techniques.
@@ 4.3.4. RNNs on FPGAs
However, long sequences can stress on-chip buffering, and performance is often bounded by external memory bandwidth or host I/O if the pipeline cannot keep data resident (Williams et al., 2009; Venieris and Bouganis, 2016). In system deployments, additional overhead can come from host-device transfers and from how sequences are batched/packed; therefore, practical FPGA RNN acceleration often co-designs the data ingress/egress pipeline with the compute datapath so that the recurrent loop remains streaming rather than I/O-bound (Nurvitadhi et al., 2017; Williams et al., 2009).
@@ 4.4.1. Different Transformer and LLM model variants
  • Encoder-only Transformers: used for representation learning (e.g., classification/embedding). They process the full sequence in parallel, so throughput is driven by batched matmuls and attention over the full context (e.g., Bidirectional Encoder Representations from Transformers (BERT)-style encoders) (Vaswani et al., 2017; Devlin et al., 2019).
@@
  • Decoder-only LLMs: used for autoregressive generation (e.g., Generative Pre-trained Transformer (GPT)-style and LLaMA-style). Inference alternates between a prefill phase (processing the prompt) and a decode phase (generating tokens). Decode is often bandwidth-limited because each token requires reading/writing KV-cache and executing relatively small GEMMs (Brown et al., 2020; Touvron et al., 2023; Kwon et al., 2023; Williams et al., 2009).
@@ 4.4.2. Key kernels and bottlenecks: attention, KV-cache, and fusion
Figure 19. Critical kernels and system bottlenecks in LLM acceleration, detailing the interplay between dense GEMM projections, bandwidth-intensive attention mechanisms, and the impact of sequence length on memory capacity.
@@
Recent acceleration strategies for Transformer/LLM inference increasingly target decode efficiency and memory traffic per token. Examples include faster IO-aware attention implementations beyond the original FlashAttention design (Dao et al., 2022; Dao, 2023), attention variants that reduce KV-cache bandwidth (e.g., grouped-query attention, which reduces the number of KV heads relative to Q heads) (Ainslie et al., 2023), and serving-time techniques such as speculative decoding that trade extra parallel work for lower end-to-end latency (Leviathan et al., 2023). Post-training quantization has also evolved into practical, high-accuracy pipelines for LLMs—including SmoothQuant, GPTQ, and activation-aware weight quantization (AWQ)—which reduce bandwidth and enable lower-precision kernels when supported by hardware and libraries (Xiao et al., 2023; Frantar and Alistarh, 2023; Lin et al., 2024; Jacob et al., 2018; Gong et al., 2024). More radically, the ”1-bit LLM” paradigm (e.g., BitNet b1.58) proposes ternary weights $\{-1,0,1\}$ to drastically reduce memory footprint and replace expensive FP16 multiplications with integer addition, promising significant energy savings if hardware support matures (Ma et al., 2024).
@@ 4.4.3. Transformers/LLMs on GPUs
When MoE is used, GPU efficiency depends on routing implementation, expert parallelism, and load balance; otherwise, conditional execution can reduce utilization and increase tail latency even when compute capacity is high (Shazeer et al., 2017; Gale et al., 2019). Overall, GPUs remain the default platform because the software ecosystem (libraries, compilers, serving runtimes) evolves rapidly with model variants and can incorporate new attention kernels and scheduling policies without changing hardware (Chen et al., 2018; Tillet et al., 2019). Continued architectural scaling is evident in recent generations such as NVIDIA’s Blackwell platform, which integrates narrower precision support (FP4/FP6) and massive chip-to-chip bandwidth to support trillion-parameter training and inference (Huang, 2024). Similarly, AMD’s Instinct MI300X series targets the same high-bandwidth, high-capacity regime by integrating CPU and GPU cores with large shared HBM pools to reduce host-device transfer overheads (AMD, 2024).
@@ 4.4.4. Transformers/LLMs on TPUs/NPUs
For large-scale training, TPU-class systems demonstrate that system co-design (HBM bandwidth, on-chip SRAM, and high-bandwidth interconnect) matters as much as per-chip compute. Scaling LLM training requires sharding and collective communication that interact with compilation decisions, and overall efficiency depends on overlap between compute and communication as well as on memory footprint (activations, optimizer state) (Jouppi et al., 2021; Micikevicius et al., 2018). Recent iterations, such as Google’s TPU v5p, further optimize this balance by increasing inter-chip interconnect bandwidth and HBM capacity specifically to handle the communication-intensive training of large generative models (Spiridonov and Ji, 2024). Reduced precision (BF16/FP16) is widely used to increase throughput and manage memory traffic (Micikevicius et al., 2018).
@@ 4.4.6. Transformers/LLMs on FPGAs
FPGA acceleration(Fig.20) for Transformers focuses on building streaming pipelines for projections/MLP and (when feasible) attention/softmax, using custom precision and on-chip buffering to reduce off-chip traffic (Zhang et al., 2024a; Nurvitadhi et al., 2017; Boutros et al., 2024). The FPGA approach mirrors classical CNN FPGA acceleration: exploit spatial pipelining and determinism, tailor datapath precision, and fuse adjacent stages so that intermediate tensors do not spill to DRAM (Venieris and Bouganis, 2016; Williams et al., 2009).
Figure 20. FPGA acceleration pipeline for Transformers, demonstrating how streaming dataflows, custom precision, and fused operators are used to handle attention mechanisms and reduce latency.
@@
System integration is a major determinant of end-to-end benefit. Host I/O (e.g., PCIe), data marshaling, and batching policies can dominate latency unless the serving pipeline is engineered so that the FPGA remains compute-fed (Nurvitadhi et al., 2017). Consequently, FPGA designs often target fixed model shapes and low-latency regimes where deterministic pipelines and tight I/O integration offset lower peak compute, while toolchain support and memory-system design remain central to achieving robust speedups (Venieris and Bouganis, 2016; Nurvitadhi et al., 2017).
@@
Beyond mainstream GPU/TPU deployments, the 2023–2025 landscape also includes a growing set of alternative datacenter accelerators (e.g., AMD Instinct MI300-class systems, Intel Gaudi-class accelerators, and AWS Trainium-class instances) (Wikipedia contributors, 2025a, c, b; Silvano et al., 2025). On the research side, there is active work on transformer-specific accelerators and system co-design: dynamic sparsity/activation pruning in transformer blocks (e.g., AccelTran) (Wang et al., 2023), 3D heterogeneous manycore organizations for end-to-end transformer execution (e.g., HeTraX) (Zhang et al., 2024b), FPGA toolflow-driven LLM inference (e.g., HLS-based transformer implementations) (Li et al., 2024a), and heterogeneous processing-in-memory (PIM) approaches that couple SRAM/HBM compute to reduce bandwidth bottlenecks (Kim et al., 2025). These efforts reinforce the same overarching lesson: as models scale and contexts grow, end-to-end speedups come from jointly optimizing kernels, memory systems, and runtimes—not from compute throughput alone (Williams et al., 2009; Kwon et al., 2023).
@@ 6. Conclusion
This survey has reviewed how neural-network acceleration has progressed from simply increasing peak compute to optimizing end-to-end execution under memory, communication, energy, and software constraints. Across model families and deployment settings, the dominant costs increasingly come from moving and managing data (activations, weights, and state) rather than from arithmetic alone. As a result, the most successful accelerator platforms are those that treat compute, memory hierarchy, interconnect, and the software stack as a single co-designed system.
We organized the hardware landscape into several major architectural families, each reflecting a different point in the design space. GPUs provide broad programmability and remain central for training and general-purpose inference, while TPU/NPU-style designs emphasize dense tensor throughput and efficiency via structured datapaths. FPGAs occupy a niche where determinism, custom precision, and streaming pipelines are valuable, and ASIC inference engines pursue maximum performance-per-watt at the cost of flexibility. Emerging architectures—including LPU-like designs for predictable token generation, in-/near-memory and analog approaches aimed at reducing data movement, and neuromorphic/event-driven processors for sparse workloads—highlight continued experimentation as models and serving requirements evolve.
Several optimization levers recur throughout the survey: precision reduction and quantization to improve effective bandwidth and energy efficiency; sparsity and pruning to reduce compute and memory footprint (when supported end-to-end by kernels and scheduling); and compilation, fusion, and layout transformations that reduce intermediate materialization and kernel overheads. For LLM serving in particular, system performance is often governed by memory capacity and bandwidth, KV-cache management, and request scheduling, making runtime policies as important as kernel-level efficiency.
Looking ahead, progress will likely be driven by three system-level needs. First, scalable LLM serving must sustain high concurrency and long contexts without becoming bandwidth- or capacity-limited. Second, dynamic and conditional computation (e.g., routed experts, tool-augmented pipelines, and heterogeneous operators) requires robust support for irregular control flow and load balance. Third, evaluation practice must mature: meaningful comparisons should report end-to-end metrics and configurations that reflect real deployment constraints, rather than relying on peak throughput alone.
In conclusion, the next generation of neural acceleration will be defined less by raw compute density and more by efficient data orchestration and hardware–software co-design. As workloads diversify and deployment expands from datacenters to edge devices, architectures that combine adaptable software stacks with well-balanced memory and communication subsystems will remain the most impactful.
Acknowledgements
This work was partially funded by DARPA (AMP, N6600120C4020; FIRE, P000050426), the NSF (FDT-Biotech, 2436801), and the Helmsley Charitable Trust (2-SRA-2017-503-M-B).
Appendix A Acronyms
Table 1. List of Acronyms.
Acronym Definition
ADC Analog-to-Digital Converter
ANN Artificial Neural Network
ASIC Application-Specific Integrated Circuit
BERT Bidirectional Encoder Representations from Transformers
BF16 Brain Floating Point (16-bit)
BNN Binarized Neural Network
BRAM Block RAM
CMOS Complementary Metal-Oxide-Semiconductor
CNN Convolutional Neural Network
DAC Digital-to-Analog Converter
DRAM Dynamic Random Access Memory
DSL Domain-Specific Language
DSP Digital Signal Processor
ECC Error Correction Code
EDP Energy-Delay Product
FLOPs Floating Point Operations
FP16 16-bit Floating Point
FP32 32-bit Floating Point
FPGA Field-Programmable Gate Array
GEMM General Matrix Multiplication
GNN Graph Neural Network
GPT Generative Pre-trained Transformer
GPU Graphics Processing Unit
GRU Gated Recurrent Unit
HBM High-Bandwidth Memory
HLS High-Level Synthesis
INT4 4-bit Integer
INT8 8-bit Integer
IO Input/Output
ISA Instruction Set Architecture
ITL Inter-Token Latency
JIT Just-In-Time
KV-cache Key-Value cache
LLM Large Language Model
LPU Language Processing Unit
LSTM Long Short-Term Memory
LUT Look-Up Table
MAC Multiply-Accumulate
MIG Multi-Instance GPU
MLP Multilayer Perceptron
MoE Mixture-of-Experts
NIC Network Interface Card
NoC Network-on-Chip
NPU Neural Processing Unit
OS Operating System
PCM Phase-Change Memory
PHY Physical Layer
QKV Query-Key-Value
ReRAM Resistive RAM
RNN Recurrent Neural Network
RTL Register Transfer Level
SIMD Single Instruction, Multiple Data
SIMT Single Instruction, Multiple Threads
SLO Service Level Objective
SM Streaming Multiprocessor
SoC System-on-Chip
SRAM Static Random Access Memory
SSD Solid State Drive
STDP Spike-Timing-Dependent Plasticity
TOPS Tera Operations Per Second
TPU Tensor Processing Unit
TTFT Time-To-First-Token